OpenStudy (anonymous):

how to do this ???? find the x. 3(x+1)+2=-7

OpenStudy (jamesj):

So I would first subtract 2 from each side. if you do that, what do you get?

OpenStudy (anonymous):

which side could u draw it pls.......

OpenStudy (jamesj):

from both sides 3(x+1) + 2 - 2 = - 7 - 2

OpenStudy (jamesj):

so what do you get when you simplify that

OpenStudy (jamesj):

3(x+1) + 2 - 2 = - 7 - 2 3(x+1) + 0 = -9 that is 3(x+1) = -9 following so far?

OpenStudy (jamesj):

still there?

OpenStudy (jamesj):

ok. good luck then.